# Rethinking Bimanual Robotic Manipulation: Learning with Decoupled Interaction Framework

### v1

@@ 1 Introduction
Figure 2: Comparisons of our Decoupled Interaction Framework with integrated control frameworks. Integrated control frameworks (a) mainly use a single model that takes the observations and states of both arms as inputs and directly outputs their actions. Our Decoupled Interaction Framework (b) first assigns an independent model to each arm to solely handle the inputs of the current arm (the yellow lines). Then, different from the naive interaction modeling in integrated control frameworks, a selective interaction module is proposed which learns its weights from its own arm to perform explicit modeling (the green and blue lines) on the exchanged state features (the green and blue dashed lines).
@@ 2 Revisiting Bimanual Robotic Manipulation
Robotic manipulation aims to utilize extensive observation-action pairs from expert demonstrations to acquire human-like skills. Previous research mainly focuses on single-arm manipulation, where the modeling of policies is primarily divided into reinforcement learning [10, 35, 6, 33, 39, 24], imitation learning [32, 29, 36, 23, 1, 5, 3, 37, 4, 18, 25, 12] and Vision-Language-Action (VLA) models [17, 42, 7, 14]. Concretely, imitation learning methods mainly focus on regression-based [32, 29, 36, 23, 1, 5] and generation-based methods [3, 37, 4, 18, 25, 12], which lay the foundation for policy designs of bimanual manipulation.
@@
$\begin{split}A_{ic}&=\pi_{ic}(O_{pcd1},O_{pcd2},S_{arm1},S_{arm2}),\\ A_{ic}&\in\mathbb{R}^{1\times 14},O_{pcd}\in\mathbb{R}^{N\times 3},S_{arm}\in\mathbb{R}^{1\times 7}.\end{split}$ (1)
@@ 3.2 The Proposed Decoupled Design Contributes to Uncoordinated Tasks
To address the problem of high-dimensional action spaces, in this paper, we propose to decouple the joint action space of both arms to reduce the space the neural network needs to explore, which helps our framework learn uncoordinated tasks more effectively. Specifically, our framework assigns an independent model to each arm, where it takes the observation and state of the current arm as inputs and outputs actions for the current arm. As shown in Fig. 1, by incorporating the decoupled design, our framework outperforms the integrated control baseline in uncoordinated tasks. Moreover, as illustrated in Fig. 1, our framework can effectively handle basic tasks like single-arm grasping, which is important in uncoordinated tasks. In conclusion, we consider that the decoupled design is beneficial for uncoordinated tasks. The loss of our decoupled design is:
@@ 3.4 A Decoupled Interaction Framework
Implementation Details. In this paper, we adopt the same point cloud backbone as DP3 [37]. Following the setup in RoboTwin [27], we utilize joint angles for proprioception and predict joint angles as actions. In the loss function, the weights for the left arm and right arm are both set to 1. Our framework is implemented using PyTorch and trained on a single NVIDIA RTX 4090 GPU for 3000 epochs with the AdamW optimizer and a batch size of 120. Additional implementation details are provided in Appendix E.
@@ 4.1 Experiment Setups
Simulation Benchmark. In this paper, we evaluate our framework on seven distinct manipulation tasks from the RoboTwin benchmark [27]. Based on the taxonomy of bimanual manipulation [15], we define two tasks as coordinated tasks and five tasks as uncoordinated tasks. Each task is designed to assess specific aspects and detailed task descriptions can be found in Appendix F.
@@
Baselines. We select several representative approaches as our baselines. For regression-based integrated control approaches, we choose ACT [40] as a baseline, which takes multi-view images as inputs and directly regresses actions of both arms with an action chunking mechanism. For generation-based integrated control approaches, we choose and modify DP [3], which takes multi-view images as inputs and directly generates actions of both arms with receding horizon control [26], and DP3 [37], which takes 3D point clouds as inputs and directly generates actions like DP, as baselines. For decoupled approaches, we refer to the architecture of Voxact-b [20] and construct a decoupled baseline that takes point clouds as inputs and outputs the actions of each arm separately. More detailed implementations of all methods can be seen in Appendix E.
@@ 4.4 Model Analysis
It should be noted that in DP3 [37], the performance of Flow Matching [19] implemented with DPM Solver++ [22] decreases compared to DDIM [30]. The explanation provided in the DP3 paper is that the high dimensions of the tasks pose a challenge to the learning of DPM Solver++. For integrated control frameworks, as seen in the first and the third lines of Tab. 4, the high dimensions of the tasks result in worse performance for PFM compared to DP3. From the second and fourth lines of Tab. 4, it can be seen that our decoupled design effectively reduces the action space that Flow Matching needs to explore, thereby fully leveraging the generative abilities of Flow Matching.
@@ 5 Conclusion
Appendix A Video Demo
A video demo is provided for both simulation and real-world manipulation experiments using our Decoupled Interaction Framework, show the effectiveness, extensibility, and practical applicability of our framework. In simulation experiments, we first collect 50 demonstrations for various tasks in the Sapien simulation environment. We then train our framework separately for each task. For real-world experiments, we utilize the Cobot Magic robotic arm and the RealSense L515 camera, employing single-view, third-person point clouds for model training and inference. Watch the video for more details. Enjoy and have fun !
@@ Appendix A Video Demo
Appendix B Analysis of Selective Interaction Module
@@
Appendix C Visualization of the Generated Manipulation Trajectories in RoboTwin
@@
Appendix D Visualization of the Generated Manipulation Trajectories in Real World
@@ Appendix F Simulation Tasks
Figure 6: Visualization of the generated manipulation trajectories of our framework in RoboTwin. We visualize different coordinated and uncoordinated tasks within various scenes. Zoom in for the best view.

### v2

@@ 1 Introduction
Figure 2: Comparisons of our Decoupled Interaction Framework with integrated control frameworks. Integrated control frameworks (a) mainly use a single model that takes the observations and states of both arms as inputs and directly outputs their actions. Our Decoupled Interaction Framework (b) first assigns an independent model to each arm to solely handle the inputs of the current arm (the yellow lines). Then, different from the naive interaction modeling in integrated control frameworks, a selective interaction module is proposed that learns its weights from its own arm to perform explicit modeling (the green and blue lines) on the exchanged state features (the green and blue dashed lines).
@@ 2 Revisiting Bimanual Robotic Manipulation
Robotic manipulation, aiming to use extensive observation-action pairs from expert demonstrations to acquire human-like skills, is underpinned by grasp detection [13, 35, 36, 39, 34, 2]. Previous research mainly focuses on single-arm manipulation, where the modeling of policies is primarily divided into reinforcement learning [11, 41, 7, 38, 45, 26], imitation learning [37, 31, 42, 25, 1, 6, 4, 43, 47, 5, 20, 27, 14] and Vision-Language-Action (VLA) models [19, 50, 8, 16, 48]. Concretely, imitation learning methods mainly focus on regression-based [37, 31, 42, 25, 1, 6] and generation-based methods [4, 43, 5, 20, 27, 14], which lay the foundation for policy designs of bimanual manipulation.
@@
$\begin{split}A_{cent}&=\pi_{cent}(O_{pcd1},O_{pcd2},S_{arm1},S_{arm2}),\\ A_{cent}&\in\mathbb{R}^{1\times 14},O_{pcd}\in\mathbb{R}^{N\times 3},S_{arm}\in\mathbb{R}^{1\times 7}.\end{split}$ (1)
@@ 3.2 The Proposed Decoupled Design Contributes to Uncoordinated Tasks
To address the problem of high-dimensional action spaces, in this paper, we propose to decouple the joint action space of both arms to reduce the space the neural network needs to explore, which helps our framework learn uncoordinated tasks more effectively. Specifically, our framework assigns an independent model to each arm, where it takes the observations and states of the current arm as inputs and outputs actions for the current arm. As shown in Fig. 1, by incorporating the decoupled design, our framework outperforms the integrated control baseline in uncoordinated tasks. Moreover, as illustrated in Fig. 1, our framework can effectively handle basic tasks like single-arm grasping, which is important in uncoordinated tasks. In conclusion, we consider that the decoupled design is beneficial for uncoordinated tasks. The loss of our decoupled design is:
@@ 3.4 A Decoupled Interaction Framework
Implementation Details. In this paper, we adopt the same point cloud backbone as DP3 [43]. Following the setup in RoboTwin [29], we utilize joint angles for proprioception and predict joint angles as actions. In the loss function, the weights for the left arm and right arm are both set to 1. Our framework is implemented using PyTorch and trained on a single NVIDIA RTX 4090 GPU for 3000 epochs with the AdamW optimizer and a batch size of 120. Additional implementation details are provided in Appendix 5.
@@ 4.1 Experiment Setups
Simulation Benchmark. In this paper, we evaluate our framework on seven distinct manipulation tasks from the RoboTwin benchmark [29]. Based on the taxonomy of bimanual manipulation [17], we define two tasks as coordinated tasks and five tasks as uncoordinated tasks. Each task is designed to assess specific aspects and detailed task descriptions can be found in Appendix 6.
@@
Baselines. We select several representative approaches as our baselines. For regression-based integrated control approaches, we choose ACT [46] as a baseline, which takes multi-view images as inputs and directly regresses actions of both arms with an action chunking mechanism. For generation-based integrated control approaches, we choose and modify DP [4], which takes multi-view images as inputs and directly generates actions of both arms with receding horizon control [28], and DP3 [43], which takes 3D point clouds as inputs and directly generates actions like DP, as baselines. For decoupled approaches, we refer to the architecture of Voxact-b [22] and construct a decoupled baseline that takes point clouds as inputs and outputs the actions of each arm separately. More detailed implementations of all methods can be seen in Appendix 5.
@@ 4.4 Model Analysis
It should be noted that in DP3 [43], the performance of Flow Matching [21] implemented with DPM Solver++ [24] decreases compared to DDIM [32]. The explanation provided in the DP3 paper is that the high dimensions of the task pose a challenge to the learning of DPM Solver++. For integrated control frameworks, as seen in the first and the third lines of Tab. 4, the high dimensions of the tasks result in worse performance for PFM compared to DP3. From the second and fourth lines of Tab. 4, it can be seen that our decoupled design effectively reduces the action space that Flow Matching needs to explore, thereby fully leveraging the generative abilities of Flow Matching.
@@ 5 Conclusion
Acknowledgments. This work was supported partially by NSFC(92470202, U21A20471), Guangdong NSF Project (No. 2023B1515040025), Guangdong Key Research and Development Program(No.2024B0101040004).
@@
1 Video Demo
A video demo is provided for both simulation and real-world manipulation experiments using our Decoupled Interaction Framework, showing the effectiveness, extensibility, and practical applicability of our framework. In simulation experiments, we first collect 50 demonstrations for various tasks in the Sapien simulation environment. We then train our framework separately for each task. For real-world experiments, we utilize the Cobot Magic robotic arm and the RealSense L515 camera, employing single-view, third-person point clouds for model training and inference. Watch the video for more details. Enjoy and have Fun!
@@ 1 Video Demo
2 Analysis of Selective Interaction Module
@@
3 Visualization of the Generated Manipulation Trajectories in RoboTwin
@@
4 Visualization of the Generated Manipulation Trajectories in Real World
@@ 6 Simulation Tasks
Figure 6: Visualization of the generated manipulation trajectories of our framework in RoboTwin. We visualize different coordinated and uncoordinated tasks within various scenes. Zoom in for best view.